OpenStudy (anonymous):

A motorist traveling at 15m/s encounters a deer in the road 43m ahead. If the maximum acceleration the vehicle's brakes are capable of is -6m/s/s, what is the maximum reaction time of the motorist that will allow him to avoid hitting the deer? Answer in units of s.

OpenStudy (ksaimouli):

ok first use v^2=vi^2+2ax

OpenStudy (ksaimouli):

find initial speed

OpenStudy (anonymous):

I haven't learned that formula yet, so I shouldn't have to use it...

OpenStudy (ksaimouli):

what did u learned

OpenStudy (anonymous):

\[v=d/t, a=(v _{f}-v _{o})/t, v(average)=(v _{o}+v _{f})/2, d=(1/2)at ^{2}, and d _{f}=d _{o}+v _{o}t+(1/2)at ^{2}\]

OpenStudy (ksaimouli):

use 4th one

OpenStudy (ksaimouli):

d=1/2

OpenStudy (ksaimouli):

.....

OpenStudy (anonymous):

v=u+at...put v=0,u=15...withacc.=-6...

OpenStudy (anonymous):

Should that get me the answer then?

OpenStudy (anonymous):

ya it will give u min. time when vechicle stops...put that in eq. s=ut+1/2at^2 u will get distance...

OpenStudy (anonymous):

I don't understand what you just said, but I used the first equation thing you gave me, plugged in the stuff you told me to and got the wrong answer

OpenStudy (anonymous):

@ksaimouli I tried what you told me to do and also got the wrong answer

OpenStudy (anonymous):

use eq. s=ut+1/2at^2...put s=43,u=15,a=-6

OpenStudy (ksaimouli):

ok i am giving the solution

OpenStudy (anonymous):

@pkjha3105 I've tried that

OpenStudy (anonymous):

try...u may get ans. easily by dis methrd..

OpenStudy (anonymous):

I ALREADY DID TRY

OpenStudy (ksaimouli):